# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients

### v1

@@ 3.6 Computational Complexity
The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (3) of Algorithm 3 necessitates a complexity of $\mathcal{O}\left(N(\mathcal{R}_{\text{ranks}}r_{1})^{2}\right)$ The residual computation in line (3) is of $\mathcal{O}({r_{1}}^{3})$ (i.e., independent of $N$) due to the matrix multiplications, with an additional $\mathcal{O}({r_{1}}^{2})$ for the Frobenius norm.
We note that henceforth $r$ will represent the maximum of $r_{\kappa}$ and $\mathcal{R}_{\text{ranks}}r_{1}$. We conclude that the overall algorithm scales as $\mathcal{{O}}(Nr^{2}+r^{3})$, which motivates keeping the overall rank $r$ as low as possible. These scalings will be verified later in this study.
3.7 Extension to High-order Temporal Discretization
The DIRK method features a stage-by-stage backward Euler-like implicit solve, with explicit evaluation of the terms on the right side of the previous Runge-Kutta stages [10]. In Appendix C.1, the mathematical formulation of the DIRK schemes is elaborated leading to the following residual equation for the $k$-th RK stage,
Expanding the operator $\mathscr{L}$ in the residual equation above results in the following $k$-th stage generalized Sylvester equation:
For the DIRK scheme, we broadly adopt the strategy outlined in [10] with minor modifications to improve efficiency. In particular, we: (i) construct the basis using the time-step of the first stage and retain it for subsequent stages, (ii) evolve the projected system for the $S_{1}$ matrix across the DIRK stages, and (iii) truncate and evaluate the residual to determine whether to admit the solution or return to step (i) to further augment the basis. Key modifications vs. [10] include truncating the solution before the residual computation, and only checking the residual at the end of the DIRK step. The reduced matrix is solved similarly to the backward Euler method outlined above, using preconditioned GMRES, where the preconditioner is based on an approximate operator, $\tilde{\mathscr{L}}$. Further details on the discretization of the advection-diffusion equation using the DIRK scheme are provided in Appendix C. Specifically, Appendix C.2 describes the adaptive-rank approach for high-order integration, Appendix C.3 extends the integration to multi-rank coefficients with DIRK schemes and presents a pseudo-algorithm.
@@ 4 Numerical Experiments
We demonstrate the computational complexity estimates and the overall efficacy of the proposed algorithm for advection-diffusion equations with several numerical examples. We assume homogeneous Dirichlet boundary conditions for all the tests for simplicity. All simulations were performed using MATLAB. For benchmarking purposes, we compare against full-rank algorithm using the Kronecker formulation inverted with MATLAB’s backslash command. Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
@@
where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively. We set algorithmic tolerances as follows. We first choose the residual tolerance, $\epsilon_{\texttt{tol}}$. One strategy to prevent rapid growth of the extended Krylov subspaces, especially when the matrix equations contain many terms, is to choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\geq\epsilon_{\texttt{tol}}$ to truncate basis vectors that are nearly linearly dependent. However, over-truncation may risk removing important information, slowing down convergence. The GMRES tolerance $\epsilon_{\texttt{GMRES}}$ is chosen as $\epsilon_{\texttt{GMRES}}<10^{-2}\epsilon_{\texttt{tol}}$ to ensure that the error in the inner loop does not pollute the outer loop. The SVD truncation threshold $\epsilon$ are set to $\epsilon<10^{-2}\epsilon_{\texttt{tol}}$. Since truncation occurs before residual evaluations, the error introduced by solution truncation must remain below the residual tolerance. Otherwise, the computed solution may meet the residual threshold, but over-truncation could prevent this acceptance.
Example 4.1: Advection-diffusion with a rank-1 advection coefficient and a rank-3 diffusion coefficient
Consider Equation (3) with a rank-1 variable advection coefficient:
$\sigma^{x}(x,y)$ $=\sigma^{1}_{1}(x)\sigma^{2}_{1}(y),$ $\sigma^{1}_{1}(x)$ $=1-x^{2},$ $\sigma^{2}_{1}(y)$ $=2y,$
$\sigma^{y}(x,y)$ $=\sigma^{1}_{2}(x)\sigma^{2}_{2}(y),$ $\sigma^{1}_{2}(x)$ $=-2x,$ $\sigma^{2}_{2}(y)$ $=1-y^{2},$
@@
$\phi^{x}(x,y)=\phi^{y}(x,y)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y),$
${\phi}^{1}_{1}(x)$ $=\exp\left(-\left(x-0.3\sin(x)\right)^{2}\right),$ ${\phi}^{2}_{1}(y)$ $=\exp\left(-\left(y-0.3\cos(y)\right)^{2}\right),$
${\phi}^{1}_{2}(x)$ $=\exp\left(-\left(x-0.6\sin(\pi x)\right)^{2}\right),$ ${\phi}^{2}_{2}(y)$ $=\exp\left(-\left(y-0.6\sin(\pi y)\right)^{2}\right),$
${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right).$
@@
$f(x,y,0)=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}\right)\right)+0.8\exp\left(-400\left((x-0.65)^{2}+(y-0.5)^{2}\right)\right).$
@@
We use a spatial grid with $N_{x}=N_{y}=300$, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $280$ to $750$. We test our adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 3 and 3 in the Appendix. The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3. The basis truncation threshold is $\epsilon_{\kappa}=10^{-3}$. The GMRES convergence threshold is $\epsilon_{\texttt{GMRES}}=10^{-8}$, and the SVD truncation threshold is $\epsilon=10^{-8}$.
@@
Figure 2 illustrates the $L_{1}$ error norm vs time stepping size characterized by $\lambda_{D}$ by showcasing the performance of the adaptive-rank integrator for BE, DIRK2, and DIRK3. The expected temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed.
@@
Figure 2: Example 4.1. Temporal convergence study of BE, DIRK2, and DIRK3 schemes.
We performed a computational complexity study to demonstrate the proposed algorithm has a linear complexity with respect to $N$, the number of grid points per dimension. The mesh size $N=N_{x}=N_{y}$ varied between $100$ and $31000$ and the time step was fixed at $\Delta t=10^{-3}$. This range of mesh sizes is chosen to prevent the $O(r^{3})$ contribution of the coefficient matrix GMRES solver from polluting the mesh-resolution scaling.
Figure 3(a) displays $N$ vs. computational time. We observe that the complexity remains linear with respect to $N$.
Figure 3: Example 4.1.(a) Computational complexity of DIRK2 adaptive-rank integrator. (b) Log-log plot of GMRES-ACS solver. The execution times were recorded using MATLAB’s timeit function.
Figure 3(b) shows the computational time required for solving the projected generalized Sylvester equation via preconditioned GMRES-ACS vs. rank for a single time step using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$. To examine the computational scaling corresponding to the expansion of the Krylov basis, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration. This figure verifies the expected cubic scaling, e.g. $\mathcal{O}(r^{3})$. From the above study, we verify that the predominant complexity of our approach is $\mathcal{O}(Nr^{2}+r^{3})$, effectively reducing the complexity of full rank methods, yet maintain their high-order convergence.
@@
We investigate next the impact of the proposed preconditioner in the S-step by comparing the performance of the GMRES solver with and without the preconditioner. Figure 4 showcases the GMRES residual history as a function of the number of iterations for both preconditioned and unpreconditioned cases for mesh sizes ranging from $200$ to $2500$, in a single time step with $\Delta t=0.01$. It is observed that preconditioning significantly accelerates convergence, with the convergence rate independent of the mesh size, which is not the case for unpreconditioned GMRES.
@@
Figure 4: Example 4.1. Semi-log plot of GMRES residuals versus iterations with and without preconditioning. The simulation is conducted for a single time step with $\Delta t=0.01$, on meshes of size $N_{x}=N_{y}=200$ in (a), $N_{x}=N_{y}=1000$ in (b), and $N_{x}=N_{y}=2500$ in (c).
Example 4.2: Adaptive Rank Evolution with Manufactured Solution
We assess next the performance of the proposed method to capture the rank adaptively using a manufactured solution defined as:
$f(x,y,t)$ $=\left(\sin(\pi x)\sin(\pi y)\right)\exp\left[-0.1\pi(t-2)^{2}\right]$
$\quad+\left(\sin(2\pi x)\sin(2\pi y)\right)\exp\left[-0.5\pi(t-10)^{2}\right]$
$\quad+\left(e^{\left[-100(x-0.3)^{2}-100(y-0.3)^{2}\right]}+e^{\left[-100(x+0.3)^{2}-100(y+0.3)^{2}\right]}\right)e^{\left[-0.5\pi(t-15)^{2}\right]}.$
This solution is initially rank-9 and fluctuates in rank over time, with significant changes at $t=1$, $10$, and $15$. Based on this solution, we analytically compute the source term $S(x,y,t)$. The diffusion coefficients are set as previously described in Example 4.1. We use a spatial grid of $N=1000$, the BE integrator and simulate until $T_{f}=18$. The relative residual tolerance is $\epsilon_{\texttt{tol}}=10^{-8}$. The GMRES convergence threshold is set to $\epsilon_{\texttt{GMRES}}=10^{-12}$. The singular-value truncation threshold is set to $\epsilon=10^{-5}$. The Krylov basis truncation threshold is $\epsilon_{\kappa}=10^{-5}$. We use a time step $\Delta t=\lambda_{D}\Delta x^{2}$ with $\lambda_{D}=1300$, or equivalently, $\Delta t=\lambda_{A}\Delta x$ with $\lambda_{A}=17$.
Figure 5(a) tracks the adaptive rank in time with our proposed adaptive rank integrator (in dashed red) vs the full-rank integrator (in solid blue). Agreement is excellent between the two. Figure 5(b) illustrates the evolution of the extended Krylov subspace over time, before and after applying the truncation step in Algorithm 3. Our approach effectively filters out low-relevance basis vectors, maximizing computational efficiency. The extended Krylov basis-augmentation history is shown in the semi-transparent red bars in Figure 5(b) (corresponding to right y-axis). It can be seen that the algorithm performs several ($\sim 3$) iterations of basis augmentation around the $t=1$, $10$, and $15$, to find an appropriate subspace to resolve the rank growth induced by the source term, settling into $\sim 1$ iteration otherwise. This demonstrates the effectiveness of the Krylov subspace formulation proposed here to find suitable low-rank subspaces as allowed by the dynamics of the solution.
Figure 5: Example 4.2. (a) Rank evolution of the adaptive-rank method using the backward Euler temporal update with a rank-3 manufactured solution. (b) Left axis: evolution of the extended Krylov subspace size over time, both before and after truncation. Right axis: number of Krylov basis iterations required for solution convergence.
Example 4.3: Swirling Deformation Flow
We consider solving the following advection-diffusion equation:
$f_{t}-\left[(1-x^{2})\cdot 2y\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot f\right]_{y}=\nu\Delta f,\quad x,y\in[-1,1]$ (35)
The initial condition is a bi-Gaussian function given by:
$f(x,y,0)=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}\right]\right)+0.8\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.65)^{2}+(y-0.5)^{2}\right]\right),$
where $\sigma=0.15$. We choose the diffusion coefficient $\nu=10^{-3}$, small enough to weaken the diffusion process. To satisfy the cell-Reynolds number (to preserve monotonicity while keeping the spatial discretization linear; see Appendix B), we set $\Delta x=9.52\times 10^{-4}$, resulting in a mesh with $N_{x}=N_{y}=2100$. This test showcases the capability of our proposed low-rank integrator to handle fine meshes, ensuring that the numerical solution adheres to the required monotonicity criteria and avoids unphysical behavior. The residual tolerance is set to $\epsilon_{\texttt{tol}}=10^{-5}$, the basis truncation threshold is set to $\epsilon_{\kappa}=10^{-3}$, the SVD truncation threshold is set to $\epsilon=10^{-5}$, and the GMRES acceptance threshold is set to $\epsilon_{\texttt{GMRES}}=10^{-10}$.
@@
Figure 6: Example 4.3. Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=7$ on a $2100\times 2100$. (a) Evolution of rank over time. (b)–(e) Solution profiles at $t=0$, $t=2.8487$, $t=5.7088$, and $t=9.9962$. The final simulation time is $T_{f}=10$.
Figure 6 demonstrates our algorithm’s ability to capture the solution rank for this challenging case. The solution rotates according to the prescribed advection coefficients and diffuses slowly over time. As the solution structure becomes more complex due to deformation, the rank increases to accurately represent its physical features. Notably, the algorithm dynamically adjusts the solution rank, which does not exceed $r_{1}=20$, two orders of magnitude smaller than the grid dimension $N=2100$.
Additionally, we evaluate the algorithm’s effectiveness in dynamically capturing the solution rank, even with large time steps. To this end, we use the proposed adaptive-rank algorithm with DIRK3 temporal integration and employ subsequent larger time steps, $\Delta t$, corresponding to CFL numbers $\lambda_{A}=1$ and $\lambda_{A}=15$. All other parameters are kept consistent with the simulation depicted in Figure 6, specifically $\epsilon_{\kappa}=10^{-3}$, $\epsilon_{\texttt{tol}}=10^{-5}$, $\epsilon=10^{-5}$, and $\epsilon_{\texttt{GMRES}}=10^{-10}$. Figure 7 shows the evolution of the generalized Krylov subspace over time, both before and after truncation, highlighting significant reductions in basis size and thus algorithmic complexity. The figure also demonstrates that the algorithm effectively captures the solution rank evolution even with fairly large (but dynamically relevant) advective CFLs at comparable performance. This demonstrates the algorithm’s capability to construct an appropriate Krylov basis, accurately evaluate the effects of the advection and diffusion operators, and dynamically truncate the subspace as needed.
Figure 7: Example 4.3. Evolution of the Extended Krylov subspace over time before and after truncation for different $\lambda_{A}$ values using DIRK3: (a) and (b) correspond to $\lambda_{A}=1$, while (c) and (d) correspond to $\lambda_{A}=15$.
Example 4.4: Constructed Steady-State Solution
@@
In this example, we construct diffusion and advection coefficients that balance each other, resulting in a non-trivial steady-state solution. Specifically, we seek an equilibrium solution satisfying the steady-state condition:
$0=\nabla\cdot\left(\mathbf{\Phi}\cdot\nabla f_{\text{eq}}-\mathbf{\Sigma}f_{\text{eq}}\right),$ (36)
@@
$\mathbf{\Sigma}=\mathbf{\Phi}\cdot\nabla\log(f_{\text{eq}}).$
This equation defines the steady-state flow from a provided diffusion coefficient and a steady-state solution. We choose the steady-state solution in the domain $0\leq x,y\leq 1$ to be:

### v2

@@ 3.6 Computational Complexity
The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)-(4) of Algorithm 4 features a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)$. The residual computation in line (4) is of $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)$ due to the matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)$ for the Frobenius norm.
It follows that, as advertised, the overall complexity of the algorithm scales as $Nr^{2}$ and as $r^{d+1}$, with $d$ the dimensionality, which will be numerically verified in Section 4.
@@ 4 Numerical Experiments
We demonstrate the overall efficacy of the proposed algorithm with several 2D and 3D numerical examples. We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (.e.g, periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (10). All simulations are performed using MATLAB. For efficient computational handling in 3D, we employ tools from the Tensor Toolbox [3] and the Htucker package [25]. Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
@@
where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively. These numbers are measures of the condition number of the associated GSEs. For all simulations presented in this section, we did not need to set a maximum number of outer xKrylov iterations because the algorithm consistently converged within a small number of iterations.
Example 4.1: 3D Advection-Diffusion
We consider a rank-1 variable advection coefficient in a 3D domain $x,y,z\in[-1,1]$:
$\sigma^{x}(x,y,z)$ $=\sigma^{1}_{1}(x)\sigma^{2}_{1}(y)\sigma^{3}_{1}(z),$ $\sigma^{1}_{1}(x)$ $=1-x^{2},$ $\sigma^{2}_{1}(y)$ $=2y,$ $\sigma^{3}_{1}(z)$ $=-2z$
$\sigma^{y}(x,y,z)$ $=\sigma^{1}_{2}(x)\sigma^{2}_{2}(y)\sigma^{3}_{2}(z),$ $\sigma^{1}_{2}(x)$ $=-2x,$ $\sigma^{2}_{2}(y)$ $=1-y^{2},$ $\sigma^{3}_{2}(z)$ $=2z$
$\sigma^{z}(x,y,z)$ $=\sigma^{1}_{3}(x)\sigma^{2}_{3}(y)\sigma^{3}_{3}(z),$ $\sigma^{1}_{2}(x)$ $=4x,$ $\sigma^{2}_{2}(y)$ $=2y,$ $\sigma^{3}_{3}(z)$ $=1-z^{2}.$
@@
$\phi^{x}(x,y,z)=\phi^{y}(x,y,z)=\sum_{i=1}^{3}\phi^{1}_{i}(x)\phi^{2}_{i}(y)\phi^{3}_{i}(z),$
${\phi}^{1}_{1}(x)$ $=\exp\left(-\left(x-0.3\sin(x)\right)^{2}\right),$ ${\phi}^{2}_{1}(y)$ $=\exp\left(-\left(y-0.3\cos(y)\right)^{2}\right),$ ${\phi}^{3}_{1}(z)$ $=\exp\left(-\left(y-0.3\sin(z)\right)^{2}\right),$
${\phi}^{1}_{2}(x)$ $=\exp\left(-\left(x-0.6\sin(\pi x)\right)^{2}\right),$ ${\phi}^{2}_{2}(y)$ $=\exp\left(-\left(y-0.6\sin(\pi y)\right)^{2}\right),$ ${\phi}^{3}_{2}(z)$ $=\exp\left(-\left(y-0.6\sin(\pi z)\right)^{2}\right),$
${\phi}^{1}_{3}(x)$ $=\exp\left(-\left(x-0.6\sin(2\pi x)\right)^{2}\right),$ ${\phi}^{2}_{3}(y)$ $=\exp\left(-\left(y-0.6\sin(2\pi y)\right)^{2}\right)$ ${\phi}^{3}_{3}(z)$ $=\exp\left(-\left(y-0.6\sin(2\pi z)\right)^{2}\right).$
@@
$f(x,y,z,t=0)$ $=0.5\exp\left(-400\left((x-0.3)^{2}+(y-0.35)^{2}+(z-0.2)^{2}\right)\right)$
$+0.8\exp\left(-400\left((x-0.65)^{2}+(y-0.5)^{2}+(z-0.55)^{2}\right)\right).$
@@
We use a spatial grid with $N=300$ in all dimensions, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $140$ to $1125$. We test our adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 3 and 3 in the Appendix. The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3, the basis truncation threshold is $\epsilon_{\kappa}=10^{-6}$, the SVD truncation threshold is $\epsilon=10^{-6}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-6}$.
We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the proposed ACS preconditioner, and (4) verification of the overall computational complexity.
@@
High-Order Temporal Convergence: Figure 1(a) illustrates the $L_{1}$ error norm vs time stepping size (characterized by $\lambda_{D}$) by showcasing the performance of the adaptive-rank integrators for BE, DIRK2, and DIRK3. The expected asymptotic temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed. Additionally, the algorithm adeptly captures the rank evolution of the solution in the x, y, and z directions, as illustrated in Figure 1(b).
@@
Figure 1: Example 4.1. (a) Temporal convergence study of BE, DIRK2, and DIRK3 schemes. (b) Rank evolution of the solution in the $x$, $y$, and $z$ directions over time for $\lambda_{D}=140$.
Effectiveness of the xKrylov Iterations: We evaluate the xKrylov subspace dimension required to achieve tolerances of $\epsilon_{\texttt{tol}}=10^{-4},10^{-5},10^{-6}$. We perform a single time-step integration with a fixed step size $\Delta t=10^{-3}$ with grid refinements in $N$. To isolate the effect of $\lambda_{D}$ (a measure of the GSE condition number) on the convergence behavior of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$. Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying spatial resolutions $N$. Our results suggest that, for a fixed tolerance, the convergence of the Krylov subspaces remains largely insensitive to grid refinement across a broad range of $\lambda_{D}$. However, for extremely fine mesh resolutions, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{1/4}$, consistently with the $\kappa^{1/4}$ scaling (with $\kappa$ the condition number) previously reported for the Poisson equation [24, 4].
Figure 2: Example 4.1. Convergence study using the xKrylov subspace method with DIRK3 adaptive-rank integrator for several outer basis tolerances, $\epsilon_{tol}$.
@@
Performance of ACS-Preconditioner: We compare the performance of the GMRES solver with and without the ACS preconditioner for the inner solve. Figure 3 illustrates the GMRES residual history over iterations for ACS-preconditioned and unpreconditioned cases. Results at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation are shown, comparing different mesh sizes. The results confirm that ACS-preconditioned GMRES delivers mesh-independent and super-linear convergence rates, which is not the case for unpreconditioned GMRES.
@@
Figure 3: Example 4.1. Log-log plots of GMRES residuals versus iterations, with and without preconditioning. Simulations are conducted for a single time step with $\Delta t=0.01$ at the third outer-basis augmentation, using grid resolutions $N_{x}=N_{y}=N_{z}=200$ in (a), $1000$ in (b), and $2500$ in (c).
Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the proposed xKrylov subspaces increases across 12 outer iterations. For these results, the grid size is fixed at $N=1000$ in 3D, and a single time step of $\Delta t=0.01$ is used. Figures 4 (b) - (d) show the GMRES residual histories for the ACS-preconditioned and unpreconditioned cases, corresponding to the second, fourth, and twelfth outer xKrylov iterations. These correspond to xKrylov subspace dimensions of $r=37$, $r=73$, and $r=217$, respectively. As the dimension of the xKrylov subspace grows, unpreconditioned GMRES convergence for the projected problem becomes progressively slower, as expected. In contrast, the ACS-preconditioner effectively maintains rapid convergence rates independent of subspace size $r$.
Figure 4: Example 4.1. (a) Average condition number for Galerkin-projected advection and diffusion operators. (b)-(d) Log-log plots of GMRES residuals versus iterations, with and without preconditioning for various subspace sizes $r_{\kappa}=37$, $r_{\kappa}=73$, and $r_{\kappa}=217$, respectively. These cases correspond to the highlighted blue data points in (a). Simulations are conducted for a single time step with $\Delta t=0.01$ and a grid of $N_{x}=N_{y}=N_{z}=1000$.
Verification of the Computational Complexity scaling: We verify the $\mathcal{O}(N)$ computational complexity claim next. We vary $N$ from $100$ to $30000$ and the time step is fixed at $\Delta t=10^{-3}$. Such range of mesh sizes is chosen to prevent the $O(r^{4})$ contributions of the ACS-GMRES solver from polluting the mesh-resolution scaling. Figure 5(a) displays $N$ vs. computational time for $10$ time steps of integration. We observe that the complexity remains linear with respect to $N$. The finite y-intercept corresponds to the $\mathcal{O}(r^{4})$ cost of the GMRES-ACS solve in 3D. Figure 5(b) displays the computational time required for the preconditioned GMRES-ACS solver vs. rank for a single time step, using BE with $N=10,000$ and $\epsilon_{\texttt{GMRES}}=10^{-10}$. For this test, we allowed the basis to grow without truncation, recording the simulation time required for the GMRES solve at each basis augmentation iteration. The expected $\mathcal{O}(r^{4})$ scaling in 3D is found. These studies verify that the complexity scaling of our approach is $\mathcal{O}(Nr^{2}+r^{4})$ in 3D.
Figure 5: Example 4.1. (a) Demonstration of ${\cal O}(N)$ computational complexity of DIRK3 adaptive-rank integrator in 3D. (b) Demonstration of ${\cal O}(r^{4})$ computational complexity of the GMRES-ACS solver in 3D. The execution times were recorded using MATLAB’s timeit function.
Example 4.2: 3D Swirling Deformation Flow
We consider next a 3D swirling deformation flow:
$f_{t}+\left[(1-x^{2})\cdot 2y\cdot-2z\cdot f\right]_{x}+\left[2x\cdot(1-y^{2})\cdot 2z\cdot f\right]_{y}+\left[-2x\cdot 2y\cdot 2(1-z^{2})\cdot f\right]_{z}=\nu\Delta f,\quad x,y,z\in[-1,1].$ (35)
We set the diffusion coefficient $\nu=1/500$. The initial condition is a bi-Gaussian function given by:
$f(x,y,0)$ $=0.5\exp\left(-\frac{1}{2\sigma^{2}}\left[(x-0.3)^{2}+(y-0.35)^{2}+(z-0.3)^{2}\right]\right)$
$+0.8\exp\left(-\frac{1}{2\sigma^{2}}\left[(x+0.5)^{2}+(y+0.5)^{2}+(z+0.5)^{2}\right]\right),$
where $\sigma=0.15$. We choose $N_{x}=N_{y}=N_{z}=1000$. The outer basis residual tolerance is $\epsilon_{\texttt{tol}}=10^{-5}$, the basis truncation threshold is $\epsilon_{\kappa}=10^{-5}$, the SVD truncation threshold is $\epsilon=10^{-5}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-10}$.
This example demonstrates the algorithm’s ability to adapt to dynamic rank evolution. The algorithm performs nine xKrylov augmentations at the initial time step, two augmentations at the third and seventh time steps, and a single augmentation otherwise. Figure 6 shows rank evolution. The solution rank rises sharply in the $x$ and $y$ directions [Figures 6 (a)-(b)]. The solution rank in the $z$ direction evolves more gently [Figure 6 (c)] and is roughly one-third of that in the $x$ and $y$ directions. Figures 6 (d)-(e) showcase how the solution deforms according to the prescribed advection coefficients and diffuses slowly over time. Figure 7 shows snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$. As the flow-driven deformation increases the complexity of the solution structure, the rank grows to capture these features; when the diffusion process starts to dominate, the rank decreases as the solution smooths. This example shows that the low-rank assumption is not intrinsic to our proposed algorithm. Even when the rank becomes moderately large, the algorithm remains capable of capturing the underlying dynamics effectively. Naturally, the benefit of a low-rank approximation decreases as the solution approaches a full-rank structure.
@@
Figure 6: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(c) Evolution of rank, xKrylov subspace before and after truncation, and number of xKrylov iterations over time for the $x$, $y$, and $z$ directions, respectively. (d)–(e) Solution profiles at $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$. The final simulation time is $T_{f}=10$.
Figure 7: Example 4.2. 3D Swirling deformation problem using DIRK3 with time step corresponding to $\lambda_{A}=14$ on a $1000\times 1000\times 1000$ grid. (a)–(f) snapshots of the solution in the $x\text{--}z$ plane at $y=-0.5$, corresponding to times $t=0$, $t=0.17167$, $t=1.7482$, $t=3.1496$, $t=5.2518$, and $t=7.0035$, respectively.
Example 4.3: A 2D Steady-State Balanced-Advection-Diffusion Problem
@@
The last example prescribes diffusion and advection coefficients that exactly balance each other, resulting in a non-trivial steady-state solution. We seek an equilibrium solution satisfying the steady-state condition:
$0=\nabla\cdot\left(\bm{\mathit{\Phi}}\cdot\nabla f_{\text{eq}}-\bm{\mathit{\Sigma}}f_{\text{eq}}\right),$ (36)
@@
$\bm{\mathit{\Sigma}}=\bm{\mathit{\Phi}}\cdot\nabla\log(|f_{\text{eq}}|).$
This equation defines the steady-state advective flow $\bm{\mathit{\Sigma}}$ for given diffusion coefficient $\bm{\mathit{\Phi}}$ and steady-state solution $f_{\text{eq}}$. We construct a steady-state solution in the domain $0\leq x,y\leq 1$ to be (up to an arbitrary constant):